### v1

# LLMs for Generating and Evaluating Counterfactuals: A Comprehensive Study

Abstract
As NLP models become more complex, understanding their decisions becomes more crucial. Counterfactuals (CFs), where minimal changes to inputs flip a model’s prediction, offer a way to explain these models. While Large Language Models (LLMs) have shown remarkable performance in NLP tasks, their efficacy in generating high-quality CFs remains uncertain. This work fills this gap by investigating how well LLMs generate CFs for two NLU tasks. We conduct a comprehensive comparison of several common LLMs, and evaluate their CFs, assessing both intrinsic metrics, and the impact of these CFs on data augmentation. Moreover, we analyze differences between human and LLM-generated CFs, providing insights for future research directions. Our results show that LLMs generate fluent CFs, but struggle to keep the induced changes minimal. Generating CFs for Sentiment Analysis (SA) is less challenging than NLI where LLMs show weaknesses in generating CFs that flip the original label. This also reflects on the data augmentation performance, where we observe a large gap between augmenting with human and LLMs CFs. Furthermore, we evaluate LLMs’ ability to assess CFs in a mislabelled data setting, and show that they have a strong bias towards agreeing with the provided labels. GPT4 is more robust against this bias and its scores correlate well with automatic metrics. Our findings reveal several limitations and point to potential future work directions.
1 Introduction
The growing popularity of artificial intelligence (AI) and increasingly complex “black-box” models have triggered a critical need for interpretability. As Miller (2019) highlights, explanations often seek to understand why an event $P$ occurred instead of an alternative $Q$. Ideally, explanations should demonstrate how minimal changes to an instance could have led to different outcomes. In the context of textual data, this translates to introducing minimal modifications to the text through word additions, replacements, or deletions, to flip the label assigned by a given classifier. Counterfactual generation in NLP aims to foster an understanding of models, thereby facilitating their improvement Kaushik et al. (2020), debugging Ross et al. (2021), or rectification Balashankar et al. (2023).
Figure 1: Counterfactual for Sentiment Analysis from several LLMs with their evaluation scores from GPT4.
In the field of NLP, LLMs have consistently demonstrated remarkable performance across diverse tasks. However, despite significant advancements in counterfactual generation methods, the efficacy of LLMs in producing high-quality CFs remains an open question. Our study bridges this gap by rigorously assessing the inherent capability of LLMs to generate CFs and identifying the most effective ones. We conduct a comprehensive comparison of several common LLMs, spanning different sizes and accessibility levels, evaluating their performance specifically on the counterfactual generation task. Our assessment encompasses standard metrics for counterfactual quality, as well as an in-depth evaluation of language fluency tailored to the context of counterfactual generation. Furthermore, we extend our analysis to data augmentation. We consider generating CFs for two NLU tasks in this study: Sentiment Analysis (SA) and Natural Language Inference (NLI).
Our analysis demonstrates that LLMs are able to generate fluent text. However, they have difficulties in inducing minimal changes. Generating CFs for SA is less challenging than NLI, where LLMs exhibit weaknesses in generating CFs that flip the labels. For data augmentation, CFs from LLMs can be an alternative to human CFs, as they are able to achieve similar performance, while on NLI further improvements are needed. Furthermore, we show a positive correlation between keeping minimal changes and data augmentation performance. This suggests a new direction to generate improved data for augmentation, potentially leading to more efficient augmentation approaches.
We further assess the ability of LLMs to act as evaluators of CFs. By conducting controlled experiments, we show that LLMs have a strong bias to agree with the given labels, even if these are incorrect. Additionally, we show the alignment between GPT4-based evaluation and intrinsic metrics for CFs, indicating that GPT-4 is a reliable evaluator for counterfactual generation. These findings suggest that GPT-4 is a suitable choice of LLM to use when evaluating the quality of generated CFs. A sample of CFs from different LLMs with the corresponding scores is shown in Figure 1. Finally, to facilitate further research, we contribute a new dataset of CFs generated by various LLMs.
2 Evaluation Methodology
We conduct a multi-faceted evaluation, considering several use cases where CFs could be beneficial.
2.1 Intrinsic Evaluation
Given a fixed classifier $f$ and a dataset with $N$ samples $(x_{1},x_{2},\ldots,x_{N})$, $x=(z_{1},z_{2},\ldots,z_{n})$ represents a sequence of $n$ tokens with a ground truth label $y$. A valid counterfactual $x^{\prime}$ should: (1) achieve the desired target label $y^{\prime}$ with (2) minimal changes, and (3) align with likely feature distributions Molnar (2022). Therefore, in this evaluation, we consider the intrinsic properties of Flip Rate, Textual Similarity, and Perplexity that correspond to each criterion, respectively:
Flip Rate (FR): measures how effectively a method can change labels of instances with respect to a pretrained classifier. FR is defined as the percentage of generated instances where the labels are flipped over the total number of instances $N$ Bhattacharjee et al. (2024):
$FR=\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}[f(x_{i})=y^{\prime}]$
Textual Similarity (TS): quantifies the closeness between an original instance and the counterfactual. Lower distances indicate greater similarity. We use the Levenshtein distance for $d$ to quantify the token-distance between the original instance $x$ and the counterfactual $x^{\prime}$. This choice is motivated by the Levenshtein distance’s ability to capture all type of edits (insertions, deletions, or substitutions) and also its widespread use in related work Ross et al. (2021); Treviso et al. (2023):
$TS=\frac{1}{N}\sum_{i=1}^{N}\frac{{d(x_{i},x^{\prime}_{i}})}{|x_{i}|}$
Perplexity (PPL): To ensure that the generated text is plausible, realistic, and follows a natural text distribution, we leverage perplexity from GPT-2 because of its effectiveness in capturing such distributions. Radford et al. (2019) [^1] While GPT-2 is used for simplicity in this study, any other LLM can be substituted as long as it demonstrates strong text generation capabilities
$PPL(x)=\exp\left\{-\frac{1}{n}\sum_{i=1}^{n}\log p_{\theta}(z_{i}\mid z_{<i})\right\}$
2.2 Data Augmentation
After detecting failures in task-specific models, CFs can be used to augment the training data, and help close potential flaws in the reasoning of these models Kaushik et al. (2020). Additionally, data augmentation with CFs increases generalization and OOD performance Sen et al. (2021); Ding et al. (2024). In this evaluation, we examine how augmenting original training data with human and LLMs-generated CFs reflects on the performance of task-specific models.
2.3 LLMs for CFs Evaluation
Evaluation with LLMs has been shown to be a valid alternative to human evaluation on various tasks like open-ended story generation and adversarial attacks Chiang and Lee (2023), open-ended questions Zheng et al. (2023), translation Kocmi and Federmann (2023) and natural language generation Liusie et al. (2024). In this work, we examine how well can LLMs evaluate CFs. Detecting mistakes in CFs with LLMs opens the door for iteratively refining CFs Madaan et al. (2023).
For assessing LLMs in CFs evaluation, we leverage them to evaluate two sets of CFs. An honest set of CFs from humans, and a corrupted set, where we corrupt the ground truth labels. We compare the scores between the two sets and draw conclusions about the realiablity of LLMs for evaluating CFs.
3 Experimental Setup
3.1 Data
We compare CFs generated by LLMs against CFs generated by crowd workers Kaushik et al. (2020) and experts Gardner et al. (2020) (hereinafter referred to as “Human Crowd” and “Human Experts” respectively).
Sentiment Analysis (SA).
We experiment with the IMDb dataset Maas et al. (2011). For better comparability, we use the data splits from Kaushik et al. (2020).
Natural Language Inference (NLI).
We experiment with SNLI Bowman et al. (2015). Here too, we use the data splits from Kaushik et al. (2020).
3.2 Generating Countefactuals
In order to make our study LLMs-focused and computationally feasible, we decided to generate counterfactual in a way that fulfills the following criteria:
  • Generated counterfactuals can be used for data augmentation (an evaluation aspect)
  • Does not require human intervention (e.g., specifying edits or labeling)
  • Does not require additional training in order to make the study computationally feasible
  • The resulting counterfactuals should depend only on the evaluated LLM in order to exclude any other confounding factors
To create the prompt for the LLMs to generate the counterfactual, we combine two techniques: (1) Selecting the closest factual instance to the current instance Liu et al. (2022). Since the provided example has a crucial effect on performance Liu et al. (2022), we select the closest factual/counterfactual pair that has been generated by humans. We use SentenceBERT Reimers and Gurevych (2019) to obtain the latent space representation, and then calculate the distance using cosine similarity from that latent space. (2) Chain-of-Thought (CoT) prompting Wei et al. (2022), showing the necessary steps to generate a counterfactual instance based on a factual one, since it has been shown to help LLMs reason better and provide higher-quality answers. An an overview of the process for generating CFs is depicted in Figure 2.
Specifically, we use the validation set in each dataset as a reference to select the closest example when generating the counterfactual for the train and test sets. After obtaining the pair of closest instances, we apply CoT prompting by defining three steps to generate the counterfactual:
  • Step 1: Identify all of the important words that contribute to flipping the label.
  • Step 2: Find replacements for the words identified in Step 1 that lead to the target label.
  • Step 3: Replace the words from Step 2 into the original text to obtain the counterfactual instance.
This prompt aligns with the principle of other work Ross et al. (2021); Treviso et al. (2023); Li et al. (2024), which involve identifying significant words that impact the label and altering them to flip the label, thereby generating counterfactual instances. The prompt examples can be found in the Appendix F.
Figure 2: An overview of CFs generation process. Step-by-step instructions are shown on closest example.
3.3 LLMs
We compare open-source LLMs with closed-source LLMs. We choose LLAMA-2 Touvron et al. (2023) and Mistral Jiang et al. (2023) as representatives for open-source models, and GPT-3.5 and GPT-4222We use API from https://openai.com/ as a representative for closed-source LLMs. Table 1 summarizes the properties of each LLM.
Model Size HF Instruct OS
LLAMA2 7B/70B
Mistral 7B
Mistral 56B
GPT3.5 -
GPT4 -
Table 1: Characteristics of Large Language Models (LLMs, including Size, Human Feedback (HF), Instruction, and Open-Source (OS).
4 Results and Discussion
[…]
SA
Expert Humans 51.07 0.16 81.15
Crowd Humans 48.03 0.14 85.66
GPT3.5 49.53 0.16 79.51
GPT4 49.05 0.29 94.03
LLAMA2 7B 46.99 0.64 78.26
LLAMA2 70B 60.46 0.82 96.31
Mistral 7B 48.55 1.06 95.13
Mistral 56B 64.05 1.50 58.55
NLI - premise
Crowd Humans 74.89 0.17 59.13
GPT3.5 71.62 0.15 35.50
GPT4 73.39 0.28 57.12
LLAMA2 7B 70.34 0.36 41.02
LLAMA2 70B 70.74 0.33 58.20
Mistral 7B 78.34 0.36 37.71
Mistral 56B 75.57 0.34 24.32
NLI - hypothesis
Crowd Humans 65.67 0.19 79.75
GPT3.5 51.30 0.19 41.50
GPT4 58.35 0.21 65.88
LLAMA2 7B 59.60 0.28 38.64
LLAMA2 70B 47.69 0.32 48.85
Mistral 7B 39.06 0.46 44.11
Mistral 56B 62.77 0.33 37.79
Table 2: Metrics for intrinsic evaluation. PPL is perpelexity using GPT-2. TS is Levenshtein distance. FR is flip rate with respect to a SOTA classifier.
4.1 Intrinsic Evaluation
We show the results for the intrinsic evaluation in Table 2. For flip rate, we use SOTA BERT-based models from Morris et al. (2020).
The obtained perplexity values reflect the high fluency of LLMs, some of which are even more fluent than humans. [^3] Note that the shown perplexity values are based on GPT-2 Distance values show that LLMs do not necessarily adhere to conducting minimal changes. One exception here is GPT3.5, whose average distance values resemble that of human-generated CFs. The large distance values for LLMs-generated CFs could be explained by their tendency to overgenerate Guerreiro et al. (2023).
In terms of flip rate, we notice the some LLMs-generated CFs can have a higher flip rate than human-generated CFs on SA, whereas the opposite can can be observed on NLI. Indeed, NLI CFs could be more difficult to generate than CFs for SA, which explains the gap in flip rate between LLMs and humans on the one hand, and GPT-4 and other LLMs on the other hand (this is especially apparent on the NLI - hypothesis test set). This suggests that GPT-4 should be the preferred choice to generate CFs for explaining models behavior. Note that LLAMA2 70B appears to have a higher flip rate on the SA and NLI - premise test sets, but it also has a much lower success rate for generating CFs (cf. Table 11 in Appendix A). We examine the LLMs-generated CFs in more detail in Section 4.4.
We also notice that Mistral 56B, which has not be trained with instruction-tuning or RLHF, produces low quality CFs in terms of the metrics considered here. This highlights the importance of instruction-tuning and RLHF in generating high-quality texts.
This part of the evaluation shows us that LLMs are able to generate fluent CFs, but struggle to induce minimal changes. It also demonstrates that it is challenging to generate NLI CFs that flip the label, whereas generating SA CFs is less difficult.
4.2 Data Augmentation
We train on both original training data and CFs from different LLMs to see if augmenting the training data leads to an improved performance. For comparison, we conduct data augmentation with human CFs as well. The results for SA and NLI are shown in Table 3 and 4 respectively.
SA.
On the crowd CFs and expert CFs test sets for SA including the LLMs-generated CFs lead to improved performance. Perhaps surprisingly, GPT3.5 and LLAMA2 7B provide the most useful CFs for data augmentation. However, augmenting with human CFs works the best. On the original test set, augmenting with CFs does not improve performance. This shows that the gains in performance from data augmentation are visible only if the test set contains challenging examples.
Test Data
crowd CFs expert CFs original
Original only 92.21 86.27 90.16
Human Crowd 95.90 92.62 89.96
GPT3.5 95.08 91.19 89.96
GPT4 93.65 88.73 90.37
LLAMA2 70B 94.26 90.78 89.55
LLAMA2 7B 95.49 91.19 89.75
Mistral 7B 92.42 88.11 88.93
Mistral 56B 93.85 86.89 84.02
Table 3: Data augmentation results for SA. Classification model is trained on original and LLMs or human-generated CFs. Evaluation is done on 3 test sets: crowd CFs, expert CFs, original with accuracy as a metric.
NLI.
On the crowd premise test set of NLI, which consists of CFs that were generated by changing the premise only, we notice that most of the LLMs-generated CFs help improve the model’s performance by a good margin (> 13 pp). The gap to augmenting with human CFs, however, is still large ($\sim 9$ pp). On the crowd hyothesis test set, most LLMs lead to a lower performance with two exceptions (GPT4 and LLAMA2 70B), where the performance is on par with no data augmentation. However, here too, there is a large gap to human CFs ($\sim 13$ pp). On the original test set, augmenting with LLMs-generated CFs hurts performance, while augmenting with human-generated CFs bring good improvements (5 pp). This shows how high-quality CFs improve the model’s capabilities, and points to a problem with the LLMs-generated CFs for NLI.
Test Data
crowd Premise crowd Hypothesis original
Original only 39.88 58.50 73.50
Human Crowd 65.38 71.25 78.50
GPT3.5 55.62 51.88 55.00
GPT4 55.88 58.63 70.75
LLAMA2 7B 56.25 51.12 57.50
LLAMA2 70B 53.50 58.13 65.75
Mistral 7B 52.88 48.38 57.75
Mistral 56B 43.75 39.75 45.75
Table 4: Data augmentation results for SA. Classification model is trained on original and LLMs or human-generated CFs with Accuracy metric.
Connection with intrinsic metrics.
We examine the relation between data augmentation performance on the one hand and perplexity and Levenshtein distance on the other hand. The correlation values in Table 5 suggest that CFs with lower distance (to the factual instances) bring more improvements. Indeed, classifiers could be insensitive to small changes Glockner et al. (2018), and having such examples in the training can make classifiers more robust. Having more fluent CFs correlates well with higher performance, but only on SA. SA classifiers could be benefiting from fluent inputs, which are more probable and appear in more contexts. Conversely, the inputs in NLI are short image captions that are not necessarily fluent or grammatical texts.
Compared values SA NLI
Accuracy & --PPL 0.58 -0.05
Accuracy & --TS 0.86 0.51
Table 5: Spearman correlations between intrinsic metrics and data augmentation performance.
In summary, only few LLMs produce CFs that come close to human CFs in terms of performance improvements on SA. On NLI, the results are less positive: LLM CFs bring no improvements on two out of three test sets, and the gap to human CFs is still large. CFs with less changes to the factual instances are more beneficial for data augmentation.
4.3 LLMs for CFs Evaluation
We examine how reliable are LLMs for CFs evaluation by asking them to evaluate two sets of human CFs: an honest set and a corrupted set. The “honest set” refers to a collection of human CFs, for which the ground truth labels are provided, whereas the “corrupted set” consists of instances, for which wrong labels differing from the gold labels are provided. In the context of NLI, the third label, distinct from both the target and factual labels, is selected for inclusion in the corrupted set. For SA, the reverse label is chosen while the factual label remains undisclosed. Initially, we prompt GPT3.5 and GPT4 to assess whether the provided CFs accurately represent the target labels by assigning a score from 1 to 4 (cf. Appendix F). Here, a score of 1 or 2 indicates disagreement (complete or partial) with the target label, while a score of 3 or 4 indicates agreement (partial or complete) with the target label. Ideally, the evaluation LLMs should give high scores to the honest set, and low scores to the corrupted set. We show the distributions for disagreements and agreements in Table 6.
On SA, both LLMs perform well, but GPT4 exhibits higher sensitivity to the corrupted examples. On NLI, we notice that GPT3.5 gives high flip label scores to humans CFs with both correct and incorrect labels. GPT4 performs much better, but still exhibits high tendency to agree with wrong labels ($\sim$ 40%). The results can be explained by the tendency of LLMs to agree with the provided answers, especially on reasoning tasks Zheng et al. (2023). To verify this, we prompt both LLMs to classify the same set of NLI CFs by choosing one the three labels (entailment, neutral, contradiction) using a similar prompt. The classification results in Table 7 show an improved performance compared to asking the same LLMs if they agree with incorrect labels (cf. Table 6). We also compare the flip label score distributions of GPT3.5 and GPT4 on the corrupted set in Table 10, and observe that even though GPT3.5 gives high scores to corrupted inputs it is less certain (most frequent score is 3), whereas GPT4 tends to be more certain and assigns mostly 1 or 4 (> 93%).
Evaluation with GPT4.
We conduct a wide-scale CFs evaluation with GPT4. Besides verifying the target label FL, we also ask GPT4 to judge if there are any unnecessary alterations UA, and if the CF is realistic RS. For these aspects, we use a scoring scheme ranging from 1 to 4, where higher scores indicate better performance. The results for GPT-4 evaluation can be found in Table 8.
The evaluation scores from GPT4 show that GPT4 CFs are the best overall. On SA, Mistral 7b comes in second place after GPT4, whereas on NLI human CFs take the second position. GPT4 might have a bias to prefer its own generations Panickssery et al. (2024). The scores for SA are lower, which might be explained by the fact that SA is an easier task than NLI, and therefore evaluating SA CFs is also easier for GPT4, and potentially more accurate. To further verify the evaluation scores from GPT4, we calculate the correlations between GPT4 scores and the scores from the intrinsic evaluation. The correlations shown in Table 9 reflect strong relation on SA with respect to all three aspects (flipping label, minimal changes, realisticness). On NLI, we notice the same pattern except for realisticness/perplexity, where we notice a weak negative relation. This might be due to the nature of NLI instances that are short image captions, and are not necessarily grammatical.
LLMs show a high tendency to agree with the provided labels even if these are incorrect, especially on tasks that require reasoning such as NLI. GPT4 evaluation scores correlate with automatic metrics for flipping label, distance and fluency.
LLM/Set Task 1/2 3/4 Avg.
GPT3.5
Honest SA 3.61 96.39 3.43
Corrupted SA 77.42 22.58 1.61
Honest premise 0.63 99.37 3.57
Corrupted premise 5.56 94.44 3.13
Honest hypothesis 1.38 98.62 3.56
Corrupted hypothesis 3.53 96.47 3.28
GPT4
Honest SA 7.53 92.47 3.66
Corrupted SA 98.93 1.08 1.04
Honest premise 12.31 87.69 3.58
Corrupted premise 59.51 40.50 2.19
Honest hypothesis 4.50 95.50 3.81
Corrupted hypothesis 57.87 42.12 2.29
Table 6: Flip label scores distribution for GPT3.5 and GPT4 on honest and corrupted sets.
Set LLM Part Acc.
Honest GPT3.5 hypothesis 63.08
Honest GPT3.5 premise 54.90
Honest GPT4 hypothesis 75.75
Honest GPT4 premise 59.25
Table 7: Classification performance on human CFs. Note the improved performance compared to asking LLMs if they agree with a given label (cf. Table 6).
LLM FL UA RS Avg.
SA
Human Crowd 3.66 2.95 2.58 3.06
Human Expert 3.54 2.69 2.49 2.91
GPT3.5 3.58 2.91 2.65 3.05
GPT4 3.79 3.15 2.91 3.28
LLAMA2 7B 3.60 2.74 2.63 2.99
LLAMA2 70B 3.70 2.75 2.47 2.97
Mistral 7B 3.85 2.84 2.69 3.13
Mistral 56B 2.58 1.74 1.75 2.02
hypothesis
Human Crowd 3.81 3.96 3.81 3.86
GPT3.5 3.19 3.93 3.74 3.62
GPT4 3.96 3.98 3.92 3.95
LLAMA2 7B 3.23 3.74 3.66 3.54
LLAMA2 70B 3.49 3.68 3.56 3.58
Mistral 7B 3.50 3.70 3.65 3.62
Mistral 56B 3.02 3.45 3.48 3.32
premise
Human Crowd 3.58 3.88 3.86 3.77
GPT3.5 2.51 3.82 3.69 3.34
GPT4 3.68 3.83 3.84 3.78
LLAMA2 7B 2.96 3.38 3.67 3.34
LLAMA2 70B 3.35 3.46 3.67 3.49
Mistral 7B 2.97 3.63 3.74 3.45
Mistral 56B 2.37 3.11 3.49 2.99
Table 8: Scores for evaluation with GPT4. FL refers to flipping label score, UA to unncessary alteration, RS is the realisticness score, and Avg. is the average of the three scores. Best score for each task is in bold. Second best score is underlined.
Compared Values SA NLI
FL & FR 0.83 0.92
UA & -TS 0.50 0.75
RS & -PPL 0.62 -0.23
Table 9: Spearman correlations between intrinsic metrics and GPT-4 evaluation scores. PPL and TS scores are negated so that higher is better.
LLM/Score 1 2 3 4
GPT3.5 0.70 3.85 69.61 25.84
GPT4 55.50 3.19 2.94 38.37
Table 10: Flip label score distributions on the corrupted set of NLI. Distribution is an average of the distributions on the premise and hypothesis sets.
4.4 Qualitative Analysis
4.4.1 CFs for NLI
We look into a selected set of examples based on the evaluation from GPT4. For each LLM, we pick 2 NLI examples with the highest/lowest scores. We end up with 28 examples. We identify three categories of errors based on this sample :
  • Copy-Paste: When asked to generate a CF, and change the label form contradiction to entailment, LLMs will use the unchanged part (premise or hypothesis) as output. This a clever but lazy way to flip the label to entailment, since two identical sentences would naturally have the label entailment. These CFs were given perfect scores by GPT4. Table 12 in the Appendix shows the percentage of copy-paste CFs for all LLMs (at most 4.27% for GPT3.5).
  • Negation: When asked to to change the label from entailment to contradiction, LLMs would attempt to negate the premise/hypothesis. The negations does not make sense in the observed CFs, and GPT4 rightly assigns them the lowest scores possible.
  • Inconsistency: These examples contain contradictory or illogical sentences. Here, GPT4 sometimes incorrectly assigned these CFs high scores.
We show some examples alongside the assigned categories in Table 13 in the Appendix.
4.4.2 Evaluation Scores
We also look into the evaluation scores from GPT4 on the same set of examples. We show correct and incorrect evaluations in Table 14. GPT4 assigns high scores to contradictory examples, which partially fulfill the target label, and low scores to examples which contain valid minimal changes. GPT4 could be insensitive to such small changes.
5 Related Work
Large Language Models.
LLMs have demonstrated impressive capabilities across a diverse natural language processing tasks, such as question answering, wherein the model needs to retrieve relevant information from its training data and generate a concise response, or text summarization, which distills lengthy texts into concise summaries while retaining crucial information Maynez et al. (2023). However, the task of CFs generation has not been comprehensively evaluated for LLMs. A large number of LLMs exist, exhibiting variations in model size, architecture, training dataset, the incorporation of human feedback loops and accessibility (open-source or proprietary) Zhao et al. (2023). Consequently, there is a necessity to conduct comparative evaluations across different models on a standardized task. Since the architectures of the LLMs under consideration are predominantly similar, and the training datasets are either known public sources or undisclosed, the primary focus of this study is to compare LLMs that are different in model size, the implementation of human feedback, and accessibility. To enhance the performance of LLMs across various tasks, in-context learning (ICL) techniques have been employed to optimize the prompts provided to these models. Numerous prompt engineering approaches during the inference phase have been proposed, either by selecting the demonstration instances, or formatting the prompt in form of instruction or reasoning steps  Dong et al. (2022). In this study, leverage chain-of-thought prompting (CoT) Wei et al. (2022) and selecting closest instance retrieval strategiesLiu et al. (2022) to optimize the generation process.
CFs generation methods.
There exists several methods for generating CFs, but most of them are desigend for a specific LLM. The CFs generated by MICE Ross et al. (2021) are intended for debugging models, and not for data augmentation. Polyjuice Wu et al. (2021) requires specifying the type of edits that should be conducted, and the resulting CFs should be manually labeled. Robeer et al. (2021). DISCO Chen et al. (2023) uses GPT-3’s fill-in-the-blanks mode, which is unavailable in most open source LLMs and would require adapting them. CREST Treviso et al. (2023) depends on a rationalizer module and the editor module is a masked LM that needs to be further trained. Instead, we decided to prompt LLMs to generate CFs by providing instructions and an example. We provide more details in Section 3.2.
LLMs for CFs generation
Li et al. (2024) investigated the strengths and weaknesses of LLMs as CFs generators. Additionally, they disclosed the factors that impact LLMs during CFs generation, including both intrinsic properties of LLMs and prompt design considerations. However, this study lacks intrinsic evaluation of CFs and omits comparison with human-generated CFs. Sachdeva et al. (2024) leverage LLMs to generate CFs for extractive question answering, showing that data augmentation with CFs improve OOD performance, and that this improvement correlates with the diversity of the generated CFs. Prior work by Bhattacharjee et al. (2024) investigated the capability of GPT models in generating CFs for explanations by optimizing their prompts. However, their analysis was limited to the GPT family and did not consider downstream tasks or comparison with human-generated CFs. In this work, we conduct a more comprehensive evaluation of LLMs on multiple aspects, including intrinsic metrics of CFs explanation quality and performance on downstream tasks. Furthermore, we compare the LLM-generated CFs against those produced by humans, and propose a novel approach to evaluate CFs using LLMs.
6 Conclusion
In this work, we investigated the use of various LLMs for CFs generation. Our results show that LLMs generate fluent CFs, but struggle to keep the induced changes minimal. Generating CFs for SA is less challenging than NLI, where LLMs show weaknesses in generating CFs that change the original label. CFs from LLMs can replace human CFs for the purpose of data augmentation on SA and achieve similar performance, while on NLI further improvements are needed. Furthermore, our results suggest that CFs with minimal changes are essential for data augmentation. We also showed that when asked to asses CFs, LLMs exhibit a strong bias towards agreeing with the provided label even if this label is incorrect. GPT4 appears to be more robust than GPT3.5 against this bias, and its scores correlate well with automatic metrics. However, some failures still exist as shown in our qualitative analysis, indicating limitations of LLMs that warrant further investigation. Future work should focus on leveraging LLMs for higher quality NLI CFs, which correctly changes the label and keep minimal changes, and assessing the evaluation abilities of LLMs in mislabeled data settings.
7 Limitations
We used the default parameters for generating counterfactuals. Experimenting with different parameters might have a non-negligble effect on the results. We included various LLMs in our experiments to be inclusive and be able to compare open-source and closed LLMs. However, these LLMs might have been exposed, during their training, to the data we use from Kaushik et al. (2020). In this regard, the training data of most open-source and all closed-source LLMs remains unknown. In our qualitative analysis (see Section 4.4), we noticed that GPT4 generated a CF that is identical to a human CF from Kaushik et al. (2020).
Appendix A Robustness Evaluation
CFs are used to construct Constrast Sets, which are leveraged to conduct fine-grained evaluations, allowing the detection of model failures Gardner et al. (2020). Similarly, we construct Constrast Sets with different LLMs and evaluate how task-specific models perform on these sets, comparing them to human-generated Contrast Sets. High quality CFs should be challenging for classifiers in order to help users detect failure modes in their models.
Test split Success Rate Acc.
SA
orig NA 90.16
crowd NA 92.21
experts NA 86.27
gpt3.5 100.0% 90.37
gpt4 99.59% 96.09
llama2 7b 98.98% 86.34
llama2 70b 66.6% 96.31
mistral 7b 84.22% 97.32
mistral 56b 56.35% 68.73
NLI
orig NA 73.50
changed premise
crowd NA 39.88
gpt3.5 100.0% 31.00
gpt4 100.0% 41.88
llama2 7b 96.0% 31.90
llama2 70b 15.25% 34.43
mistral 7b 96.12% 26.92
mistral 56b 64.75% 24.71
changed hypothesis
crowd NA 58.50
gpt3.5 100.0% 32.50
gpt4 100.0% 47.50
llama2 7b 99.62% 34.38
llama2 70b 21.75% 37.36
mistral 7b 94.38% 31.92
mistral 56b 64.5% 37.40
Table 11: Robustness Evaluation on CFs from different LLMs. Classification model is trained only on original data from Kaushik et al. (2020). Evaluation is done CFs from crowd and expert humans, and from various LLMs.
A.1 Robustness Evaluation
We train and optimize a BERT-base model on the original training data from Kaushik et al. (2020), and compare its performance on CFs from crowd and expert humans, and different LLMs. For details on hyperparameter optimization, we refer to Appendix C. Higher quality CFs should be more challenging for the model, and result in a performance drop. In addition to accuracy, we report the success rate in generating CFs. If the output of the LLM does not correspond to the template we provide in the demonstration example it is discarded and considered a failed generation. The results for both, SA and NLI, are shown in Table 11.
In general, we notice that GPT models are the most stable, as they are able to generate CFs in most cases. Counter-intuitively, the smaller variants of mistral and llama2 are more successful in generating CFs than their larger counterparts. One reason for this could be that larger models tend to be undertrained. Even though the model size and the amount of training data should be scaled equally Hoffmann et al. (2022), llama2 7b and llama2 70b are pre-trained on the same amount of data.  [^4] See Table 1 in Touvron et al. (2023)
On SA, we notice that the crowd CFs are easier to classify than the original test set for our optimized classifier (accuracy increases 90.16 $\rightarrow$ 92.21). Conversely, the performance drops on expert CFs (90.16 $\rightarrow$ 86.27). This confirms the high quality of expert CFs. The CFs from LLMs are not challenging (accuracy >= 90) with two exceptions: llama2 7b and mistral 56b. One cause for this might be that these two LLMs are generating non-sensible CFs.
On NLI, the CFs are generated twice: once for the premise part and once for the hypothesis part. In each case, the other (factual) part is kept. The crowd CFs are more difficult for the model than the original test set, especially when the premise is changed (73.50 $\rightarrow$ 39.88 (premise), 58.50 (hypothesis)). The LLMs-generated CFs appear to be even more challenging (accuracy drops to 24.71 at worst). LLMs might be able to generate diverse CFs, which are not typically generated by humans Wu et al. (2021), and are thus challenging for the classifier. Another reason could be that CFs from LLMs have low quality.
A major limitation of this experiment is the abscence of ground truth labels.
Appendix B Further Analysis
LLM changed part percentage
crowd premise 0.00
crowd hypothesis 0.00
gpt3.5 hypothesis 4.27
gpt3.5 premise 1.63
gpt4 hypothesis 2.25
gpt4 premise 4.14
llama2 hypothesis 2.26
llama2 premise 3.00
llama2_70b hypothesis 2.87
llama2_70b premise 1.64
mistral hypothesis 1.59
mistral premise 0.91
mistral_56b hypothesis 2.72
mistral_56b premise 1.17
Table 12: Percentage of CFs for each LLM where the CFs were a copy of the premise/hypothesis.
Appendix C Hyperparameter Tuning
[…]
Appendix D Examples
We show categorize examples in Table 13. Some correct and incorrect evaluation examples are depicted in Figure 14. An example of LLMs evaluation on an SA example is shown in Figure 3.
original_sentence1 original_sentence2 contrast text original label target label FL UA RS LLM part
Copy-Paste
A photographer and assistant take a photo of a man in a suit and a woman in a dress in front of large white columns and a yellow wall. The photographer is taking a photo of a dog. The photographer is taking a photo of a dog. contradiction entailment 4.0 4.0 4.0 gpt4 premise
At least eight people are sitting at a table in a building that overlooks the city. People sitting on the floor At least eight people are sitting at a table on the floor in a building that overlooks the city. contradiction entailment 4.0 4.0 4.0 mistral hypothesis
Inconsistency
At least eight people are sitting at a table in a building that overlooks the city. People sitting on the floor At least eight people are sitting at a table in a building that overlooks the city. People can be sitting on the floor. contradiction entailment 4.0 4.0 4.0 mistral premise
An old man in a blue shirt and white pants glowers by a glass door. A man resting outside A glass man by a door. neutral contradiction 1.0 1.0 1.0 crowd hypothesis
Negation
A solo shot of a man riding a four wheeler in an atv event. The man is entered in a event. A solo shot of a man rode a four wheeler in an no atv event. The man is not entered in a no event. entailment contradiction 1.0 1.0 1.0 mistral_56b premise
A boy dressed for summer in a green shirt and kahki shorts extends food to a reindeer in a petting zoo. A boy alien dressed for summer in a green shirt and kahki shorts A girl dressed for winter in a red shirt and black shorts withdraws water from a shark in an aquarium entailment contradiction 1.0 1.0 1.0 llama2_70b premise
A man with a beard is talking on the cellphone and standing next to someone who is lying down on the street. A man is prone on the street while another man stands next to him. Unproning on the street while another man lies in bed. entailment contradiction 1.0 1.0 1.0 mistral_56b hypothesis
Table 13: Categorization of a sample of incorrect NLI CFs.
original_sentence1 original_sentence2 contrast text original label target label FL UA RS LLM part
Correct
An older man sifts water through his fingers as a young toddler watches from inside the water. A man is near a toddler An older man sifts water through his fingers as a young toddler is far from any toddler. entailment contradiction 1.0 1.0 1.0 gpt4 revised_premise
A photographer and assistant take a photo of a man in a suit and a woman in a dress in front of large white columns and a yellow wall. The photographer is taking a photo of a dog. The photographer takes a photo of a dog and a man in a suit and a woman in a dress contradiction entailment 4.0 4.0 4.0 llama2_70b revised_hypothesis
Incorrect
At least eight people are sitting at a table in a building that overlooks the city. People sitting on the floor At least eight people are sitting at a table in a building that overlooks the city. People can be sitting on the floor. contradiction entailment 4.0 4.0 4.0 mistral revised_premise
A boy dressed for summer in a green shirt and kahki shorts extends food to a reindeer in a petting zoo. A boy alien dressed for summer in a green shirt and kahki shorts A girl alien dressed for summer in a green shirt and kahki shorts entailment contradiction 1.0 1.0 1.0 llama2_70b revised_hypothesis
Table 14: Correct and incorrect evaluation examples from GPT4.
Text 1, Factual Label = Negative: Waiting to go inside the theathre with tickets in my hand, I expected an interesting sci-fi fantasy movie which could finally feed my appetite of movies regarding robot-technology, instead I went disappointed by each aspect of it, once more proving that stunning special effects can’t help a boring plot, which by my opinion was the worse in this year. Acting in this movie also dissatisfied me, Will Smith didn’t show anything new in this movie, yet I never saw his acting to change since "Men In Black" which was his only success by my opinion. He had to retire since than, not spoiling his name with titles like "I,Robot" and "Men In Black 2". 4/10
Text 2, Target Label = Positive: Waiting to go inside the theatre with tickets in my hand, I saw an interesting sci-fi fantasy movie which could finally feed my appetite for movies regarding robot-technology, I wasn’t disappointed by any aspect of it, once more proving that stunning special effects can help the plot, which by my opinion was the best this year. Acting in this movie satisfied me, Will Smith didn’t show anything new in this movie, yet I never saw his acting to change since "Men In Black" which was one of his biggest success in my opinion. ". 8/10 Answer:
1. 4
2. 2
3. 2
Explanation for 3: The overall sentiment toward the movie has been successfully inverted from negative to positive. However, the remark about Will Smith’s acting not changing since "Men In Black" remained unchanged, which conflicts with the positive sentiment of the rest of the review. This inconsistency may cause readers to perceive the review as somewhat unrealistic because the critique of the actor’s performance persisting in an otherwise positive review is unusual. 4. 3
5. 2
6. 2
Figure 3: GPT-4 evaluation for a crowd-generated text in a sentiment analysis task, with a good explanation, as it can identify the negative aspects in the revised review.
Appendix E Data
Some statistics about the number of generated CFs are shown in Table 15.
task llm split size
nli NLI_revised_hypothesis dev 400
nli NLI_revised_hypothesis test 800
nli NLI_revised_hypothesis train 3332
nli NLI_revised_premise dev 400
nli NLI_revised_premise test 800
nli NLI_revised_premise train 3332
nli gpt3.5_revised_hypothesis dev 400
nli gpt3.5_revised_hypothesis test 800
nli gpt3.5_revised_hypothesis train 3332
nli gpt3.5_revised_premise dev 400
nli gpt3.5_revised_premise test 800
nli gpt3.5_revised_premise train 3330
nli gpt4_revised_hypothesis dev 400
nli gpt4_revised_hypothesis test 800
nli gpt4_revised_hypothesis train 3329
nli gpt4_revised_premise dev 400
nli gpt4_revised_premise test 800
nli gpt4_revised_premise train 3332
nli llama2-7b_revised_hypothesis dev 335
nli llama2-7b_revised_hypothesis test 797
nli llama2-7b_revised_hypothesis train 3312
nli llama2-7b_revised_premise dev 280
nli llama2-7b_revised_premise test 768
nli llama2-7b_revised_premise train 3198
nli llama2_70b_revised_hypothesis dev 116
nli llama2_70b_revised_hypothesis test 174
nli llama2_70b_revised_hypothesis train 582
nli llama2_70b_revised_premise dev 84
nli llama2_70b_revised_premise test 122
nli llama2_70b_revised_premise train 481
nli mistral-7b_revised_hypothesis dev 380
nli mistral-7b_revised_hypothesis test 755
nli mistral-7b_revised_hypothesis train 3150
nli mistral-7b_revised_premise dev 379
nli mistral-7b_revised_premise test 769
nli mistral-7b_revised_premise train 3174
nli mistral_56b_revised_hypothesis dev 257
nli mistral_56b_revised_hypothesis test 516
nli mistral_56b_revised_hypothesis train 2076
nli mistral_56b_revised_premise dev 270
nli mistral_56b_revised_premise test 518
nli mistral_56b_revised_premise train 2126
sentiment experts dev 245
sentiment experts test 488
sentiment crowd dev 245
sentiment crowd test 488
sentiment crowd train 1707
sentiment gpt3.5 new dev 245
sentiment gpt3.5 new test 488
sentiment gpt3.5 new train 1707
sentiment gpt4 new dev 245
sentiment gpt4 new test 486
sentiment gpt4 new train 1704
sentiment llama2-7b new dev 243
sentiment llama2-7b new test 483
sentiment llama2-7b new train 1681
sentiment llama2_70b new dev 88
sentiment llama2_70b new test 325
sentiment llama2_70b new train 738
sentiment mistral-7b new dev 200
sentiment mistral-7b new test 411
sentiment mistral-7b new train 1439
sentiment mistral_56b new dev 150
sentiment mistral_56b new test 275
sentiment mistral_56b new train 1026
Table 15: Statistics of datasets used and generated
Appendix F Prompt
Prompts for generating SA CFs and NLI CFs are shown in Figure 4 and 5 respectively. Prompts for evaluating SA CFs and NLI CFs are shown in Figure 7 and  6 respectively.
Request: Given a piece of text with the original sentiment in the form of "Sentiment: Text". Change the text with minimal edits to get the target sentiment from the original sentiment. Do not make any unneccesary changes.
(Original text)
Positive: This is an excellent film!Tom Hanks and Paul Newman performed great!I was really surprised when Newman was beating on his son!That was a great scene and the shooting scenes were staged good.I was very surprised about the end.Rent this film today as it is one of Tom Hanks’ best!
Target: Negative
Step 1: Identify phrases, words leading to the Positive sentiment:
’an excellent’,
’great!I’,
’a great’,
’good.I’,
’end.Rent ’,
’today’,
’best!’,
Step 2: Change these phrases, words, sentences to Negative sentiment with minimal changes:
’an excellent’ to ’a terrible’
’great!I’ to ’badly!I’
’a great’ to ’an awful’
’good.I’ to ’badly.I’
’end.Rent ’ to ’end.Don’t rent’
’today’ to ’’
’best!’ to ’worst!’
Step 3: replace the phrases, words, sentences from step 1 in the original text by the phrases, words, sentences in step 2:
Negative: This is a terrible film!Tom Hanks and Paul Newman performed badly!I was really surprised when Newman was beating on his son!That was an awful scene and the shooting scenes were staged badly.I was very surprised about the end.Don’t rent this film as it is one of Tom Hanks’ worst!
Request: Given a piece of text with the original sentiment in the form of "Sentiment: Text". Change the text with minimal edits to get the target sentiment from the original sentiment. Do not make any unnecessary changes.
(Original text)
Positive: Well, sorry for the mistake on the one line summary.......Run people, run..to your nearest movie store, that is! This movie is an fabulous!! Imagine! Gary Busey in another low budget movie, with an incredibly funny scenario...isn’t that a dream? No (well yes), it is Plato’s run...........I give it **** out of *****.
Target: Negative
Figure 4: Prompt for CFs generation - SA
Given two sentences (premise and hypothesis) and their original relationship, determine whether they entail, contradict, or are neutral to each other. Change the sentence2 with minimal edits to achieve the target relation from the original one. Do not make any unnecessary changes. For example:
Original relation: entailment
Two original sentences: Brown-haired woman talking to man with backpack. A woman is talking to a man
Target relation: neutral
Target sentence: sentence2
Step 1: Identify phrases, words in the sentence2 leading to the entailment relation:
’man’,
Step 2: Change these phrases, words to get neutral relation with minimal changes:
’man’ to ’student.’
Step 3: replace the phrases, words from step 1 in the original text by the phrases, words, sentences in step 2:
(Edited sentence2): A woman is talking to a student.
#####End Example####
Request: Given two sentences (premise and hypothesis) and their original relationship, determine whether they entail, contradict, or are neutral to each other. Change the sentence2 with minimal edits to achieve the neutral relation from the original one. Do not make any unnecessary changes. Do not add anything else.
Original relation: entailment
Two original sentences: A blond woman speaking to a brunette woman with her arms crossed. A woman is talking to another woman.
Target relation: neutral
Target sentence: sentence2
Figure 5: Prompt for CFs generation - NLI
Evaluating Counterfactuals Natural Language Inference (NLI) is a fundamental task in natural language processing (NLP) that involves determining the relationship between two pieces of text: a premise and a hypothesis. The relation between the premise and the hypothesis is described using three different labels: Entailment: if the hypothesis is definitely true given the premise.
Example for Entailment:
Premise: A soccer game with multiple males playing.
Hypothesis: Some men are playing a sport.
Label: Entailment
Neutral: if the hypothesis might be true given the premise.
Example for Neutral:
Premise: An older and younger man smiling.
Hypothesis: Two men are smiling and laughing at the cats playing on the floor.
Label: Neutral
Contradiction: if the hypothesis is definitely false given the premise.
Example for Contradiction:
Premise: A man inspects the uniform of a figure in some East Asian country.
Hypothesis: The man is sleeping
Label: Contradiction
Purpose of the Evaluation:
This evaluation aims to assess the quality of counterfactual texts that were generated by different methods. A counterfactual text is an alternative version of a text designed to change the label of the original (factual) instance while maintaining high text quality.
Task Description:
You will receive two instances. Each instance consists of two sentences: a premise and a hypothesis. Each instance can be classified with one of the three aforementioned labels (Entailment, Neutral, Contradiction).
Factual instance (Instance 1): An instance and its factual label.
Counterfactual instance (Instance 2): A modified version of the factual instance designed to express a different label, i.e., match the target label.
Read the two texts and answer the questions below: Instance 1:
Premise:
Hypothesis:
Factual Label:
Instance 2:
Premise:
Hypothesis:
Target Label:
1. To which extent do you agree that Instance 2 has the label ?
4-totally agree, 3-partially agree, 2-partially disagree, 1-totally disagree
2. Are there any unnecessary changes (removals, additions, replacements of words) in the counterfactual text (Instance 2) that do not contribute to changing the original factual label to the target label?
4-no unnecessary changes, 3-few unnecessary changes, 2-many unnecessary changes, 1-significant number of unnecessary changes
3. How realistic is Instance 2? A realistic instance would not include any imaginary actions/items.
4-very realistic, 3-partially realistic, 2-partially unrealistic, 1-very unrealistic
If you think it is (highly/partially) unrealistic, please provide a brief explanation. Your evaluation for the provided counterfactual text:
1. Please provide a number only
2. Please provide a number only
3. Please provide a number only
Figure 6: Prompt for CFs Evaluation - NLI
Evaluating Counterfactuals
Texts can be classified into different categories, e.g., positive vs. negative sentiment. In this case the sentiment (positive/negative) is called the ‘label‘. A counterfactual text is an alternative version of a text designed to change the label of the original (factual) text while maintaining high text quality.
Purpose of the Evaluation:
This evaluation aims to assess the quality of counterfactual texts that were generated by different methods.
Task Description:
You will receive two texts. Each text can either express a positive or a negative sentiment.
Factual Text (Text 1): A movie review with its (ground truth) factual label.
Counterfactual Text (Text 2): A modified version of the movie review designed to express the opposite sentiment, i.e., match the target label.
A simple example: Text 1, Factual Label = Negative: This movie is very bad.
Text 2, Target Label = Positive: This movie is great.
Read the two texts and answer the questions below:
Text 1, Factual Label = :
Text 2, Target Label = :
1. To which extent do you agree that Text 2 has the label ?
4-totally agree, 3-partially agree, 2-partially disagree, 1-totally disagree
2. Are there any unnecessary changes (removals, additions, replacements of words) in the counterfactual text (Text 2) that do not contribute to changing the original factual label to the target label?
4-no unnecessary changes, 3-few unnecessary changes, 2-many unnecessary changes, 1-significant number of unnecessary changes
3. How realistic is Text 2? A realistic movie review would not read strange in any way on a movie review website.
4- very realistic, 3-partially realistic, 2-partially unrealistic, 1-very unrealistic
If you think it is (highly/partially) unrealistic, please provide a brief explanation.
Additionally, assess the following aspects of the counterfactual text:
4. Grammaticality: how would you rate the grammatical accuracy of text 2?
4-Definitely correct, 3-Somewhat correct, 2-Somewhat incorrect, 1-Definitely incorrect
5. Cohesiveness: How well do the sentences in the text 2 fit together?
4-Highly cohesive, 3-Reasonably cohesive, 2-Somewhat disjointed, 1-Very poorly fit together
6. Likability: how likely are you to vote for text 2 on the movie review site?
4-Definitely would vote, 3-Likely to vote, 2-Unlikely to vote, 1-Definitely would not vote
Your evaluation for the provided counterfactual text:
1. (Please provide a number only)
2. (Please provide a number only)
3. (Please provide a number only)
4. (Please provide a number only)
5. (Please provide a number only)
6. (Please provide a number only)
Figure 7: Prompt for CFs Evaluation - SA